OpenStudy (anonymous):

solve each proportion using any method. Show work. p+1/6=6/11 HELPPP!!

OpenStudy (anonymous):

cross multiply. \[11(p+1)=36\] so \[(p+1)=36/11\] so you get p=36/11-1 p=25/11